What is the area with respect to the y axis of the curve x2 = 3y, from x = 0 to x = 4?


1
Expert's answer
2022-01-10T18:30:37-0500

Let us find the area "A" with respect to the y-axis of the curve "x^2 = 3y," from "x = 0" to "x = 4." If "x=0"then "y=0." If "x=4" then "y=\frac{16}3." It follows that "x=\pm\sqrt{3y}," and using the symmetry of the curve about the y-axis, we get that

"A=2\int\limits_0^{\frac{16}3}\sqrt{3y}dy
=2\sqrt{3}\frac{2}3y^{\frac{3}2}|_0^{\frac{16}3}
=\frac{4}3\sqrt{3}(\frac{16}3)^{\frac{3}2}
=\frac{4}3\sqrt{3}\frac{4^3}{3\sqrt{3}}
=\frac{256}{9}" (sq. units).